# OpenGuardrails: An Open-Source Context-Aware AI Guardrails Platform

### v1

@@ Abstract
As large language models (LLMs) become increasingly integrated into real-world applications, safeguarding them from unsafe, malicious, or privacy-violating content is of critical importance. We present OpenGuardrails, the first open-source project that provides both a context-aware safety & manipulation-detection model and a deployable platform for comprehensive AI guardrails. OpenGuardrails protects against content-safety risks, model-manipulation attacks (e.g., prompt injection, jailbreaking, code-interpreter abuse and generation/execution of malicious code), and data leakage. Content-safety and model-manipulation detection are implemented by a unified large model; data-leakage identification and redaction are performed by a separate lightweight NER pipeline (e.g., presidiso-style models or regex-based detectors). The system can be deployed as a security gateway or API-based service with enterprise-grade, fully private deployment options. OpenGuardrails achieve state-of-the-art(SOTA) performance on safety benchmarks, excelling in both prompt and response classification across English, Chinese, and multilingual tasks. All models are released under the Apache 2.0 license for public use.
@@ 1 Introduction
OpenGuardrails is designed to address this full risk spectrum in a unified, deployable system. Unlike prior efforts that open-source either models or rule-based tools in isolation, OpenGuardrails provides:
@@ 3.3 Distinctive Position of OpenGuardrails
    Unified LLM architecture: Both content-safety and model-manipulation detection are handled by the same LLM, unlike multi-model pipelines [6].

### v2

# OpenGuardrails: A Configurable, Unified, and Scalable Guardrails Platform for Large Language Models

@@ Abstract
As large language models (LLMs) are increasingly integrated into real-world applications, ensuring their safety, robustness, and privacy compliance has become critical. We present OpenGuardrails, the first fully open-source platform that unifies large-model-based safety detection, manipulation defense, and deployable guardrail infrastructure. OpenGuardrails protects against three major classes of risks: (1) content-safety violations such as harmful or explicit text generation, (2) model-manipulation attacks including prompt injection, jailbreaks, and code-interpreter abuse, and (3) data leakage involving sensitive or private information. Unlike prior modular or rule-based frameworks, OpenGuardrails introduces three core innovations: (1) a Configurable Policy Adaptation mechanism that allows per-request customization of unsafe categories and sensitivity thresholds; (2) a Unified LLM-based Guard Architecture that performs both content-safety and manipulation detection within a single model; and (3) a Quantized, Scalable Model Design that compresses a 14B dense base model to 3.3B via GPTQ while preserving over 98% of benchmark accuracy. The system supports 119 languages, achieves state-of-the-art performance across multilingual safety benchmarks, and can be deployed as a secure gateway or API-based service for enterprise use. All models, datasets, and deployment scripts are released under the Apache 2.0 license.
@@ 1 Introduction
OpenGuardrails is designed to address this full risk spectrum in a unified, deployable system. As shown in Figure 1, OpenGuardrails achieves consistently higher F1 scores than existing guard models across multilingual safety benchmarks, demonstrating both superior generality and robustness. Unlike prior efforts that open-source either models or rule-based tools in isolation, OpenGuardrails provides:
@@ 3.3 Distinctive Position of OpenGuardrails
    Unified LLM architecture: Both content-safety and model-manipulation detection are handled by the same LLM, unlike multi-model pipelines (llamafirewall, ).
@@ Discussion
Model Size and Quantization Trade-off
[…]
This trade-off demonstrates that semantic capacity—rather than parameter count alone—drives safety detection quality, while modern quantization enables practical deployment without compromising reliability. As a result, OpenGuardrails achieves near-SOTA performance at a fraction of the computational cost, confirming that large-scale semantic modeling followed by quantization is an effective paradigm for real-world LLM safety systems.